OpenStudy (anonymous):

(3,?)for the equation 3x -4y=2

OpenStudy (anonymous):

you must plug in 3 into the equation wherever you see x, then evluate: 3(3)-4y=2

OpenStudy (karatechopper):

y= 7/4

OpenStudy (anonymous):

your right

OpenStudy (karatechopper):

thank yyou

OpenStudy (anonymous):

Ok i'm not getting it

OpenStudy (anonymous):

check my first post

OpenStudy (anonymous):

what you were given is a coordinate (3,?), okay so this is only the x part of the ordered pair (x,y). To get the y coordinate you need to evulaute the function at 3.

OpenStudy (anonymous):

so 3(3) whic

OpenStudy (anonymous):

3(3) which is 9

OpenStudy (karatechopper):

u plug in 3, as x, then solve the equation like this: 3x-4y=2 3(3)-4y=2 9-4y=2 (9-9)-4y=(2-9) -4y=-7 -4/-4 * y= -7/-4 y=7/4

OpenStudy (karatechopper):

tht is how i solved it, helpful?

OpenStudy (anonymous):

Whew!!!!!! thats a lot is their a simple way

OpenStudy (anonymous):

thas as simple as its gets

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

unless you graphed it out, and then saw what value the function takes on at x=3

OpenStudy (karatechopper):

ya i did tht all to make u understand, so u know.... working it out makes u know what ur doing, i learned tht the hard way